OpenStudy (anonymous):

each percent is an annual interest rate. in the formula y=a*b^x what value would you use for b? 3.7%

OpenStudy (zehanz):

3.7% means 3.7 per cent, or: 3.7/100 = 0.0037. You start with 100%, so after one year you have 103.7% = 103.7/100=1.0037. So b=1.0037 is your growth rate.